OpenStudy (starryknight):

Solve for X in the following system of equations. -6x - 5y = 3 7x + 9y = -32 =-18 =7 =-36 =2 please explain to solve the problem

OpenStudy (anonymous):

one way you can solve it without taking too long, plug in the options given below to see which on gets the correct answer listed if that makes sense. if it doesnt, ill give the answer if you want.

OpenStudy (starryknight):

how do you plug them in ?

OpenStudy (anonymous):

try each answer choice. and substitute the variables by the way y=-9

OpenStudy (starryknight):

i need help fineding X not y

OpenStudy (starryknight):

:)

OpenStudy (anonymous):

Multiply the first equation by 7: 7(-6x -5y) = 7*3 Multiply the second equation by 6: 6(7x + 9y) = - 6*32 Add up the 2 lines, you will get y.